THE LABORATORY ROTATIONAL SPECTRUM OF ISO-PROPYL CYANIDE AND AN ASTRONOMICAL SEARCH IN SAGITTARIUS B2(N)

\underline{HOLGER S. P. MÜLLER}, I. Physikalisches Institut, Universität zu Köln, 50937 Köln; and Max-Planck-Institut für Radioastronomie, 53121 Bonn, Germany; A. COUTENS, A. WALTERS, CESR, Université de Toulouse (UPS), and CNRS, 31028 Toulouse, France; J.-U. GRABOW, Institut für Physikalische Chemie und Elektrochemie, Lehrgebiet A, Universität Hannover, 30167 Hannover, Germany; A. BELLOCHE, K. M. MENTEN, Max-Planck-Institut für Radioastronomie, 53121 Bonn, Germany; S. SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany.

We have carried out a molecular line survey of Sagittarius B2(N) in the 3 mm region with selected recordings at 2 and 1.3 mm to probe the chemical complexity in massive star-forming regions. Noteworthy results include the detection of aminoacetonitrile,^{a} a possible precursor of the aminoacid glycine, the detection of ^{13}C isotopologs of vinyl cyanide,^{b} and the detection of ethyl formate as well as normal-propyl cyanide.^{c} The heavy atoms in the latter molecule form a chain. An isomer with a branched structure, iso-propyl cyanide, also exists, but its rotational spectrum has only been recorded in few transitions up to 40 GHz.^{d}^{e} Therefore, laboratory measurements were extended. The molecule is rather asymmetric (κ = −0.5766) with a strong a-dipole moment component of 4.05 (2) D and a still sizable c-component of 1.4 (2) D.^{e} Measurements in Köln were carried out in selected regions between 40 and 600 GHz. Since the c-type transitions appeared to be weaker than predicted additional Stark (and also zero-field) measurements have been carried out in Hannover between 6 and 20 GHz. We will present results of these laboratory spectroscopic investigations as well as the outcome of a search for the molecule in our Sgr B2(N) line survey.

aA. Belloche, K. M. Menten, C. Comito, H. S. P. Müller, P. Schilke, J. Ott, S. Thorwirth, C. Hieret, Astron. Astrophys. 482 (2008) 179; Erratum 492 (2008) 796.

bH. S. P. Müller, A. Belloche, K. M. Menten, C. Comito, P. Schilke, J. Mol. Spectrosc. 251 (2008) 319.

cA. Belloche, R. T. Garrod, H. S. P. Müller, K. M. Menten, C. Comito, P. Schilke, Astron. Astrophys. (2009), accepted.

dG. E. Herberich, Z. Naturforsch. 22a (1967) 543.

eJ. R. Durig, Y. S. Li, J. Mol. Struct. 21 (1974) 289.